# Training-free Camera Control for Video Generation

### v3

@@ Abstract
We propose a training-free and robust solution to offer camera movement control for off-the-shelf video diffusion models. Unlike previous work, our method does not require any supervised finetuning on camera-annotated datasets or self-supervised training via data augmentation. Instead, it can be plugged and played with most pretrained video diffusion models and generate camera controllable videos with a single image or text prompt as input. The inspiration of our work comes from the layout prior that intermediate latents hold towards generated results, thus rearranging noisy pixels in them will make output content reallocated as well. As camera move could also be seen as a kind of pixel rearrangement caused by perspective change, videos could be reorganized following specific camera motion if their noisy latents change accordingly. Established on this, we propose our method CamTrol, which enables robust camera control for video diffusion models. It is achieved by a two-stage process. First, we model image layout rearrangement through explicit camera movement in 3D point cloud space. Second, we generate videos with camera motion using layout prior of noisy latents formed by a series of rearranged images. Extensive experiments have demonstrated its superior performance in both video generation quality and camera motion alignment compared with other finetuned methods. Furthermore, we show the capacity of CamTrol generalizing to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation. Videos available at https://lifedecoder.github.io/CamTrol/.
@@ 1 Introduction
As a more appealing and content-richer modality, videos differ from images by including an extra temporal dimension. This temporal aspect provides increased versatility for depicting diverse and dynamic movements, which can be decomposed into object motion, background transitions and perspective changes. Recent years have witnessed the rapid development and splendid breakthrough of video generation with text prompt or images as input instructions (Li et al., 2023; Hong et al., 2022; Ho et al., 2022; Luo et al., 2023; Zeng et al., 2024; Blattmann et al., 2023a; Brooks et al., 2024; Ge et al., 2023; Fei et al., 2023), and demonstrated the inestimable potential of diffusion models to synthesis realistic videos. While these video generation models have made progress in generating videos with highly dynamic objects and backgrounds (Zeng et al., 2024; Blattmann et al., 2023a; Li et al., 2023), most of them fail to provide camera control for the generated videos.
@@
The difficulty of controlling camera trajectory in videos primarily arises from two aspects. The initial challenge lies in the inadequacy of annotated data. Most video annotations lack of descriptions, especially precise descriptions of video’s camera movements. As a result, video generation models trained on these data often fail to interpret text prompts related to camera motions and generate correct outputs. One solution to mitigate the data insufficiency problem is to mimic videos with camera movements through simple data augmentation (Yang et al., 2024a). However, these methods could only handle simple camera motions like zoom or truck, and have trouble in dealing with more complicated ones. The second challenge lies in the effort of additional finetuning required for controlling camera movements. As camera trajectories could be sophisticated, they sometimes cannot be accurately elaborated using naïve text prompts alone. Common solutions (Wang et al., 2023; He et al., 2024) proposed to embed camera parameters into diffusion models through learnable encoders and perform extensive finetuning on large-scale datasets with detailed camera trajectories. However, such datasets like RealEstate10k (Zhou et al., 2018) and MVImageNet (Yu et al., 2023) are intensively limited in scale and diversity due to the difficulty associated with data collection, in this way, these finetuning methods demand substantial resources but exhibit limited generalizability to other types of data. Lack of annotations and heavy finetuning effort make camera control a challenging task in video generations.
@@
In this work, we attempt to address these issues through a training-free solution to offer camera control for off-the-shelf video diffusion models. We begin by introducing two core observations underpin that video diffusion models can achieve camera movement control in a training-free manner. First, we find that base video models could produce results with rough camera moves by integrating specific camera-related text into input prompts, such as camera zooms in or camera pans right. This simple implementation, though not very accurate and always leads to static or wrong motions, shows the natural prior knowledge learnt by pretrained models about following different camera trajectories. The other observation is the effectiveness video models have exhibited in adapting to 3D generation tasks. Recent works (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023) find that leveraging pretrained video models as initialization helps drastically improve the performance of multi-view generations, demonstrating their strong ability of handling perspective change. The two crucial observations reveals the hidden power of video models for camera motion control, thus, we seek to find a way to evoke this innate ability, as it already exists in the model itself.
@@
We propose CamTrol, which offers camera control for off-the-shelf video diffusion models in a training-free but robust manner. CamTrol is inspired by the layout prior that noisy latents hold towards generation results: As pixels in noisy latents change their positions, corresponding rearrangement will also occur to the output and leads to layout modification. Considering camera moves could also be seen as a kind of layout rearrangement, this prior can serve as an efficacious hint providing video model with information of specific camera motions. Specifically, CamTrol consists of a two-stage procedure. In stage I, explicit camera movements are modeled in 3D point cloud representations and produce a series of rendered images indicating specific camera movements. In stage II, layout prior of noisy latents are utilized to guide video generations with camera movements. Compared with previous works, CamTrol require no any additional finetuning utilizing camera annotated datasets, nor does it need self-supervised training based on data augmentation. Extensive experiments have demonstrated its superior performance in both video generation quality and camera motion alignment against other finetuned methods. Furthermore, we show the capacity of CamTrol generalizing to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation.
@@ 2 Related Work
While methods aim for controlling video foundation models constantly emerge (Ma et al., 2024; Liu et al., 2023; Feng et al., 2023), there are few works explore how to manipulate camera motion of generated videos. Earlier work (Hao et al., 2018) controls motion trajectory via warping image through densified sparse flow and pixel fusion, similar ideas also appear later in Chen et al. (2023) and Yin et al. (2023). Besides utilizing optical flow, two main techniques for implementing video camera control are via self-supervised augmentation or additional finetuning. Yang et al. (2024a) disentangles object motion with camera movement and incorporates extra layers to embed camera motions, where model is trained in a self-supervised manner by augmenting input videos to stimulate simple camera movements. He et al. (2024) and Wang et al. (2023) train an additional camera encoder and integrate the output into temporal attention layers of U-Net. Guo et al. (2023) learns new motion pattern via LoRA (Hu et al., 2021) and finetuning with multiple reference videos.
@@
One of the most natural advantages of diffusion model comes from its pixel-wise noisy latents formed during denoising process. These latents hold strong causuality towards output and directly determine what the result looks like, meanwhile have robust error-resilience as they are perturbed by Gaussian noises across different scales. Numerous work have exploited the convenience of this noise prior to attain controllable generation, such as image-to-image translation (Meng et al., 2021), pixel-level manipulation (Nichol et al., 2022), image inpainting (Lugmayr et al., 2022) and semantic editing (Choi et al., 2021; Hou et al., 2024). Recent research has shown that even sampled from Gaussian distribution, the initial noise of diffusion process still have significantly influence to the layout of generated contents (Mao et al., 2023). In other work, noise prior is used to guarantee temporal consistency among video frames (Luo et al., 2023), or to trade-off between fidelity and diversity of image editing (Kim et al., 2022).
@@
Similar to how most video generation models using the groundwork laid by image foundation models (Blattmann et al., 2023b; Esser et al., 2023; Singer et al., 2022; Wu et al., 2023), training of 3D generation model also relies heavily on pretrained 2D video models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024). These methods either finetunes with rendered videos directly (Blattmann et al., 2023a; Chen et al., 2024; Melas-Kyriazi et al., 2024; Han et al., 2024), or adds camera embedding for each view as extra condition (Voleti et al., 2024; Shi et al., 2023). Video foundation models have shown to be particularly beneficial in generating consistent multi-view rendering of 3D objects, demonstrates their inherent abundant prior knowledge for handling camera pose change.
@@ 3 Training-free Camera Control for Video Generation
Figure 1: Pipeline of CamTrol. In stage I, camera movements are modeled through explicit 3D point cloud . In stage II, layout prior of noisy latents are utilized to guide video generation.
CamTrol takes two stages to evoke the innate camera control ability hidden in foundation models. In Sec. 3.1, we will describe how to model explicit camera movement for video generation. In Sec. 2, we will elaborate on video’s motion control with the guidance of noise layout prior.
@@ 3.1 Camera Motion Modeling
To evoke pretrained video diffusion model’s ability of dealing with camera perspective changes, hints of camera motion should be injected to diffusion model in a proper way. While simply concatenating camera trajectories with text prompt is incomprehensible for original model, previous works (Wang et al., 2023; He et al., 2024) introduce additional embedder to encode camera parameters and finetune with limited annotated data (Zhou et al., 2018; Yu et al., 2023), which are data-hungry yet lack of generalization ability. Other method (Yang et al., 2024a) constructs camera motions by self-supervised augmentations, but could only handle a few easy camera controls. Thus, we seek a more efficient and robust way to guide the model towards camera controllable.
Considering perspective change of video is originally caused by camera movements in 3D space, we resort to 3D representation for providing generation models with explicit motion hints. Specifically, we choose point cloud as the intermediate representation, in which space we can expediently manipulate camera poses for simulating diverse movements. One extra benefit that point cloud brings is its data-efficiency: By utilizing inpainting techniques, only one single input image is required for the whole point cloud reconstruction, this sidesteps the effort of large-scale finetuning.
@@
We start by lifting the pixels in input image plane to 3D point cloud representations. In practice, the input image can be either user-defined or created by image generators like Stable Diffusion (Rombach et al., 2022). Given an input image $\mathbf{I}_{0}\in\mathbb{R}^{3\times H\times W}$, we first estimate its depth map $\mathbf{D}_{0}$ using off-the-shelf monocular depth estimator ZeoDepth (Bhat et al., 2023). By combining image and its depth map, point cloud $\mathcal{P}_{0}$ can be initialized as:
@@
To get consistent images from multiple views, we set camera motion as a trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each including a rotation matrix and translation matrix representing camera’s pose and position. At each step $i$, we project the point cloud back to camera plane using function $\psi$ and get a rendered image with perspective change: $\mathbf{I}_{i}=\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})$. By calculating extrinsic matrices of corresponding movement, we obtain a series of camera motions including zoom, tilt, pan, pedestal, truck, roll and rotate, enabling flexible camera movements. Detailed definitions of these movements are elaborated in Appendix B. By combining basic trajectories, hybrid camera movements can be attained and produce videos with cinematic charm. What’s more, benefit from explicit camera motion modeling, our method could support trajectories with precise extrinsics, which means it can generate videos with any complicated camera motion.
@@
When perspective changes, there can be vacancies appear as some areas are unoccupied within the point cloud. To get more reasonable results, we employ image inpainting model (Rombach et al., 2022) to fill up the holes for new renderings, with a mask distinguishing the known points from nonexistent ones. This operation guarantees coherence between known views and novel views in 2D space. After inpainting, image is lifted again onto 3D space and gradually complete the whole point cloud representation. During this process, points between adjacent views may become misaligned since depth estimator only estimates relative depth, further leads to inconsistency in both 3D point cloud and rendered images. To avoid this situation, we adopt depth coefficient optimization (Chung et al., 2023) at each step of the camera movement, formed as:
@@
where $\tilde{\mathbf{I}}_{i}$ and $\tilde{\mathbf{D}}_{i}$ refer to the inpainted image and its depth map respectively, $d_{i}$ denotes depth coefficient to be optimized, and $M$ refers to the overlapping region between $\mathcal{P}_{i}$ and $\mathcal{P}_{i-1}$, as other areas are not shared for calculating $\ell_{1}$ loss.
Thus, we get a set of images refer to the input and indicate specific camera movement:
@@ 3.2 Layout Prior of Noise
Figure 2: Transition of samples between two distinct distributions. As the layout-arranged images being inverted by adding random noise, the distribution of their noisy latents will gradually converge to their video counterpart (green area), finally forming nearly Gaussian latents with specific layout change. This information is then inherited during video generation process. Steps of camera motion inversion determine the trade-off between video diversity and motion fidelity.
With camera motion modeling, we obtain a sequence $\mathbf{V}_{0}=\left\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right\}\in\mathbb{R}^{N\times 3\times H\times W}$ of rendered images adhering to a specific camera trajectory. Note that quality of rendered images are not perfect since single input image only leads to sparse point cloud reconstruction, besides, these renderings are static, thus they could not use directly as video frames. To form an ideal video, we need to find a way that satisfies three requirements: 1) camera motions should be maintained; 2) video should be encouraged with more dynamics; and 3) quality imperfection should be compensated.
@@
Recent work on diffusion models have demonstrated the strong controllability of its noisy latents (Meng et al., 2021; Mao et al., 2023), the causality and error-resilience they hold towards final output make them a convenient yet powerful tool for controllable generation of diffusion models. Particularly for initial noise, even sampled from Gaussian distribution, it still have significant influence on the layout of generated image (Mao et al., 2023), so that rearranging the noise pixels will make content in output relocate as well. For instance, if all pixels in initial noise shift to right by a certain distance, it is likely that generated output reflect a similar shift. This reminds us that the impact of camera movement on images could also be regarded as a kind of layout rearrangement, where pixels change their positions caused by viewpoint change. In a similar way, videos can be reorganized following camera motion if their noisy latents change accordingly.
Inspired by this, we first construct a series of noisy latents indicating specific camera movements. It can be intuitively done by employing diffusion’s inversion process on rendered image sequence $\mathbf{V}_{0}$. Latent at timestep $t_{0}$ can be calculated as follows, where $\bar{\alpha}_{t}$ are variances of the scheduler:
@@
Because the rendered images $\mathbf{V}_{0}$ share common pixels in certain regions, their latents also have relevance to each other in a way indicating pixels’ move. Moreover, while being perturbed with random noise, blank spaces and flawed regions in $\mathbf{V}_{0}$ can be further filled with randomness, providing video model with more possibilities to generate and correct them.
@@
After camera motion inversion, noisy latents presenting camera movements are then passed through the backward process of video diffusion model, utilizing their layout controllability to guide video generation. Leveraging prior knowledge of base video model, the generation process also bestows video with rational dynamic information. In this way, explicit camera movements are injected into video diffusion model in an appropriate and training-free fashion. Starting from noisy motion latents at timestep $t_{0}$, the generation step can be represented as:
@@
Here $\epsilon_{\theta}$ denotes the video model for noise prediction and $\sigma_{t}$ determines whether the denoising process is deterministic or probabilistic, we set $\sigma=1$ to encourage diversity of generation results.
@@
Leveraging noise prior guidance in diffusion model could lead to a trade-off between generation’s fidelity and diversity (Meng et al., 2021; Hou et al., 2024), where results that hold more faithfulness towards guidance tend to decline in generation quality. In this task, similar circumstance also occurs as model is required to be guided by some imperfect renderings while generate a reasonable and dynamic video. The key factor to balance the trade-off problem lies in the choice of $t_{0}$. When larger $t_{0}$ applied, generation bears more resemblance to original guidance $\mathbf{V}_{0}$, yet lacks of rationality and dynamics to be an appealing video. Instead, smaller $t_{0}$ leads to well-generated video, but is less aligned with desired camera motion. Empirically, we find larger $t_{0}$ works better for motions with moderate intensity, and for those with relatively drastic move, smaller $t_{0}$ shows preferable performance. The process of stage II is illustrated in Fig. 2.

### v4

@@ Abstract
We propose a training-free and robust solution to offer camera movement control for off-the-shelf video diffusion models. Unlike previous work, our method does not require any supervised finetuning on camera-annotated datasets or self-supervised training via data augmentation. Instead, it can be plug-and-play with most pretrained video diffusion models and generate camera-controllable videos with a single image or text prompt as input. The inspiration for our work comes from the layout prior that intermediate latents encode for the generated results, thus rearranging noisy pixels in them will cause the output content to relocate as well. As camera moving could also be seen as a type of pixel rearrangement caused by perspective change, videos can be reorganized following specific camera motion if their noisy latents change accordingly. Building on this, we propose CamTrol, which enables robust camera control for video diffusion models. It is achieved by a two-stage process. First, we model image layout rearrangement through explicit camera movement in 3D point cloud space. Second, we generate videos with camera motion by leveraging the layout prior of noisy latents formed by a series of rearranged images. Extensive experiments have demonstrated its superior performance in both video generation and camera motion alignment compared with other finetuned methods. Furthermore, we show the capability of CamTrol to generalize to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation. Videos available at https://lifedecoder.github.io/CamTrol/.
@@ 1 Introduction
As a more appealing and content-richer modality, videos differ from images by including an extra temporal dimension. This temporal aspect provides increased versatility for depicting diverse and dynamic movements, which can be decomposed into object motion, background transitions, and perspective changes. Recent years have witnessed the rapid development and splendid breakthroughs of video generation with text prompts or images as input instructions (Li et al., 2023; Hong et al., 2022; Ho et al., 2022; Luo et al., 2023; Zeng et al., 2024; Blattmann et al., 2023a; Brooks et al., 2024; Ge et al., 2023; Fei et al., 2023), demonstrating the immense potential of diffusion models to synthesize realistic videos. While these video generation models have made progress in generating videos with highly dynamic objects and backgrounds (Zeng et al., 2024; Blattmann et al., 2023a; Li et al., 2023), most of them fail to provide camera control for the generated videos.
@@
The difficulty of controlling camera motion in videos arises primarily from two aspects. The initial challenge lies in the inadequacy of annotated data, as most video annotations lack detailed descriptions, particularly about the camera movements. As a result, video generation models trained on these data often fail to interpret text prompts related to camera motions and generate correct outputs. One solution to mitigate the data insufficiency problem is to mimic videos with camera movements through simple data augmentation (Yang et al., 2024a). However, these methods can only handle simple camera motions like zoom or truck, and have trouble dealing with more complicated ones. The second challenge is the extra finetuning required for camera control and its inherent limitations. As camera trajectories could be sophisticated, they sometimes cannot be accurately elaborated using naïve text prompts alone. Common solutions (Wang et al., 2023; He et al., 2024) proposed to embed camera parameters into diffusion models through learnable encoders and perform extensive finetuning on large-scale datasets with detailed camera trajectories. However, such datasets as RealEstate10k (Zhou et al., 2018) and MVImageNet (Yu et al., 2023) are highly limited in scale and diversity due to the difficulty associated with data collection; in this way, these finetuning methods demand substantial resources but exhibit limited generalizability to other types of scenes. Lack of annotations and the constraints of finetuning make camera control a challenging task in video generations.
@@
In this work, we attempt to address these issues through a training-free solution to offer camera control for off-the-shelf video diffusion models. We begin by introducing two core observations underpinning the idea that video diffusion models can achieve camera movement control in a training-free manner. First, we find that base video models could produce results with rough camera moves by integrating specific camera-related text into input prompts, such as camera zooms in or camera pans right. This simple implementation, though not very accurate and always leading to static or wrong motions, shows the natural prior knowledge learned by pretrained models about following different camera trajectories. The other observation is the effectiveness of video models in adapting to 3D generation tasks. Recent works (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023) find that leveraging pretrained video models as initialization helps drastically improve the performance of multi-view generations, demonstrating their strong ability to handle perspective change. The two crucial observations reveal the hidden power of video models for camera motion control. Therefore, we seek to find a way to invoke this innate ability, as it already exists in the model itself.
@@
We propose CamTrol, which offers camera control for off-the-shelf video diffusion models in a training-free but robust manner. CamTrol is inspired by the layout prior that noisy latents possess for the generation results: As pixels in noisy latents change their positions, corresponding rearrangement will also occur in the output, leading to layout modification. Considering camera moves can also be seen as a type of layout rearrangement, this prior can serve as an effective hint, providing the video model with information about specific camera motions. Specifically, CamTrol consists of a two-stage procedure. In stage I, explicit camera movements are modeled in 3D point cloud representation and produce a series of rendered images indicating specific camera movements. In stage II, the layout prior of noisy latents is utilized to guide video generations with camera movements. Compared with previous works, CamTrol requires no additional finetuning on camera-annotated datasets, nor does it need self-supervised training based on data augmentation. Extensive experiments have demonstrated its superior performance in both video generation quality and camera motion alignment against other finetuned methods. Furthermore, we show CamTrol’s capability to generalize to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation.
@@ 2 Related Work
While methods aiming to control video foundation models continue to emerge (Ma et al., 2024; Liu et al., 2023; Feng et al., 2023), there are few works exploring how to manipulate the camera motion of generated videos. Earlier works (Hao et al., 2018) control motion trajectory via warping images through densified sparse flow and pixel fusion. Similar ideas also appear later in Chen et al. (2023) and Yin et al. (2023). Apart from utilizing optical flow, two main techniques for implementing video camera control are self-supervised augmentation or additional finetuning. Yang et al. (2024a) disentangles object motion with camera movement and incorporates extra layers to embed camera motions, where the model is trained in a self-supervised manner by augmenting input videos to simulate simple camera movements. He et al. (2024) and Wang et al. (2023) train additional camera encoders and integrate the outputs into the temporal attention layers of U-Net. Guo et al. (2023) learns new motion patterns via LoRA (Hu et al., 2021) and finetuning with multiple reference videos.
@@
One of the most natural advantages of diffusion models comes from its pixel-wise noisy latents formed during the denoising process. These latents have a strong causal effect on the output and directly determine what the result looks like, meanwhile having robust error resilience as they are perturbed by Gaussian noise across different scales. Numerous works have exploited the convenience of this noise prior to attain controllable generation, such as image-to-image translation (Meng et al., 2021), pixel-level manipulation (Nichol et al., 2022), image inpainting (Lugmayr et al., 2022) and semantic editing (Choi et al., 2021; Hou et al., 2024). Recent research has shown that, although sampled from a Gaussian distribution, the initial noise of the diffusion process still has a significant influence on the layout of generated content (Mao et al., 2023). In other works, noise prior is used to guarantee temporal consistency between video frames (Luo et al., 2023), or to trade off between fidelity and diversity of image editing (Kim et al., 2022).
@@
Similar to how most video generation models use the groundwork laid by image foundation models (Blattmann et al., 2023b; Esser et al., 2023; Singer et al., 2022; Wu et al., 2023), the training of 3D generation models also relies heavily on pretrained 2D video models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024). These methods either finetune with rendered videos directly (Blattmann et al., 2023a; Chen et al., 2024; Melas-Kyriazi et al., 2024; Han et al., 2024), or add camera embedding as an extra condition for each view (Voleti et al., 2024; Shi et al., 2023). Video foundation models have been shown to be particularly beneficial in generating consistent multi-view rendering of 3D objects, demonstrating their inherent and abundant prior knowledge for handling camera pose change.
@@ 3 Training-free Camera Control for Video Generation
Figure 1: Pipeline of CamTrol. In stage I, camera movements are modeled through explicit 3D point cloud. In stage II, the layout prior of noisy latents is utilized to guide video generation.
CamTrol takes two stages to invoke the innate camera control ability hidden in foundation models. In Sec. 3.1, we will describe how to model explicit camera movement for video generation. In Sec. 2, we will elaborate on video motion control with the guidance of noise layout prior.
@@ 3.1 Camera Motion Modeling
To invoke pretrained video diffusion models’ ability to deal with camera perspective changes, hints of camera motion should be injected into the diffusion model in a proper way. While simply concatenating camera trajectories with text prompts is incomprehensible to the original model, previous works (Wang et al., 2023; He et al., 2024) introduce additional embedders to encode camera parameters and finetune with limited annotated data (Zhou et al., 2018; Yu et al., 2023), which are data-hungry yet lack generalization ability. Other methods (Yang et al., 2024a) construct camera motions by self-supervised augmentations, but can only handle a few easy camera controls. Thus, we seek a more efficient and robust way to guide the model towards being camera-controllable.
Considering that video perspective changes are originally caused by camera movements in 3D space, we leverage 3D representation to provide generation models with explicit motion hints. Specifically, we choose point cloud as the intermediate representation, in which we can easily manipulate camera poses to simulate diverse movements. An additional benefit that point cloud brings is its data efficiency: By utilizing inpainting techniques, only one single input image is required for the entire point cloud reconstruction. This eliminates the need of large-scale finetuning.
@@
We start by lifting pixels in the input image plane into 3D point cloud space. In practice, the input image can be either user-defined or created by image generators like Stable Diffusion (Rombach et al., 2022). Given an input image $\mathbf{I}_{0}\in\mathbb{R}^{3\times H\times W}$, we first estimate its depth map $\mathbf{D}_{0}$ using the off-the-shelf monocular depth estimator ZeoDepth (Bhat et al., 2023). By combining the image and its depth map, we initialize the point cloud $\mathcal{P}_{0}$ as:
@@
To get consistent images from multiple views, we model the camera motion as a trajectory of extrinsic matrices $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each including a rotation matrix and a translation matrix representing camera pose and position. At each step $i$, we project the point cloud back to the camera plane using $\psi$ and get a rendered image with perspective change: $\mathbf{I}_{i}=\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})$. By computing the extrinsic matrices of different movements, we obtain a series of camera motions including zoom, tilt, pan, pedestal, truck, roll, and rotation, enabling flexible camera movements. Detailed definitions of these movements are provided in Appendix B. Basic trajectories are combined to produce hybrid camera movements, adding cinematic appeal to the generated videos. Furthermore, benefiting from explicit camera motion modeling, our method can support trajectories with precise extrinsics, which means it can generate videos with arbitrarily complex camera motions.
@@
When perspective changes, holes can appear as some areas are unoccupied within the point cloud. To obtain more reasonable results, we employ image inpainting (Rombach et al., 2022) to fill the gaps in new renderings, with a mask distinguishing the known points from the nonexistent ones. This operation ensures coherence between the known and novel views in 2D space. After inpainting, the image is lifted back into 3D space and gradually completes the whole point cloud. During this process, points between adjacent views may become misaligned since the depth estimator only estimates relative depth, leading to inconsistencies in 3D point cloud and rendered images. To avoid this situation, we adopt depth coefficient optimization (Chung et al., 2023) at each step of the camera movement, formed as:
@@
where $\tilde{\mathbf{I}}_{i}$ and $\tilde{\mathbf{D}}_{i}$ refer to the inpainted image and the corresponding depth map, respectively, $d_{i}$ denotes the depth coefficient to be optimized, and $M$ refers to the overlapping region between $\mathcal{P}_{i}$ and $\mathcal{P}_{i-1}$, as other areas are not shared for calculating $\ell_{1}$ loss.
Thus, we get a set of images that correspond to the input and indicate specific camera movements:
@@ 3.2 Layout Prior of Noise
Figure 2: Transition of samples between two distinct distributions. As the layout-arranged images are inverted by adding random noise, the distribution of their noisy latents will gradually converge to that of their video counterpart (green area), eventually forming a nearly Gaussian latent with specific layout change. This information is then inherited during the generation process. The inversion step determines the trade-off between video diversity and motion fidelity.
With camera motion modeling, we obtain a sequence $\mathbf{V}_{0}=\left\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right\}\in\mathbb{R}^{N\times 3\times H\times W}$ of rendered images adhering to a specific camera trajectory. Note that the quality of rendered images is not perfect since a single input image only leads to sparse point cloud reconstruction. Moreover, these renderings are static, thus they cannot be used directly as video frames. To form an ideal video, we need to find a way that satisfies three requirements: 1) camera motions should be maintained; 2) the video should be enriched with more dynamics; and 3) quality imperfection should be compensated.
@@
Recent work on diffusion models has demonstrated the strong controllability of their noisy latents (Meng et al., 2021; Mao et al., 2023), the causality and error-resilience they possess with respect to the final output make them a convenient yet powerful tool for controllable generation in diffusion models. Particularly for initial noise, even when sampled from a Gaussian distribution, it still has significant influence on the layout of the generated image (Mao et al., 2023). For instance, if all pixels in initial noise shift to the right by a certain distance, it is likely that the generated output reflects a similar shift. This reminds us that the impact of camera movement on images can also be interpreted as a type of layout rearrangement, where pixels change their positions due to viewpoint change. In a similar way, videos can be reorganized following camera motion if their noisy latents change accordingly.
Inspired by this, we first construct a series of noisy latents indicating specific camera movements. It can be done intuitively by employing diffusion’s inversion process on the rendered image sequence $\mathbf{V}_{0}$. Latent at timestep $t_{0}$ can be calculated as follows, where $\bar{\alpha}_{t}$ is the variance of the scheduler:
@@
Because the rendered images $\mathbf{V}_{0}$ share common pixels in certain regions, their latents are also correlated to each other in a way reflecting pixel movement. Moreover, by introducing random noise, blank spaces and flawed regions in $\mathbf{V}_{0}$ can be adaptively completed, providing the video model with more possibilities to generate and correct them.
@@
After camera motion inversion, noisy latents representing camera movements are then passed through the backward process of the video diffusion model, leveraging their layout controllability to guide video generation. Using prior knowledge of the base model, the generation process also infuses the video with rational dynamic information. In this way, explicit camera movements are injected into the video diffusion model in an appropriate, training-free fashion. Starting from noisy motion latents at timestep $t_{0}$, the generation step can be represented as:
@@
Here $\epsilon_{\theta}$ denotes the video model for noise prediction and $\sigma_{t}$ determines whether the denoising process is deterministic or probabilistic. We set $\sigma=1$ to encourage diversity in generation process.
@@
Leveraging noise prior guidance in diffusion models could lead to a trade-off between generation fidelity and diversity (Meng et al., 2021; Hou et al., 2024), where results that are more faithful to the guidance tend to degrade in generation quality. In this task, similar circumstances also occur, as the model is required to be guided by some imperfect renderings while generating a reasonable video. The key factor in balancing the trade-off problem lies in the choice of $t_{0}$. When a smaller $t_{0}$ is applied, the generated output bears more resemblance to original guidance $\mathbf{V}_{0}$, but lacks rationality and dynamics to be an appealing video. Instead, a larger $t_{0}$ leads to better-generated videos but makes them less aligned with the desired camera motion. Empirically, we find that a smaller $t_{0}$ works better for motions with moderate intensity, and for those with relatively drastic movements, a larger $t_{0}$ shows preferable performance. The process of stage II is illustrated in Fig. 2.